OpenStudy (anonymous):

Find the slope of the line that passes through the given points. Reduce all fractional answers to lowest terms. (-3, 4) and (-5, -6) m =

OpenStudy (anonymous):

did you do x1-x2 and y1-y2 yet

OpenStudy (the_fizicx99):

M = (y2 - y1) / (x2 - x1) M = -6-4 / -5 + 3 M = -10 / -2 M = 5